A chemist has two solutions of boric acid, one 11 percent boric acid and the other 19 percent boric acid. How much of each must he use to obtain 25 liters of a solution that is 17 percent of boric acid?
Answer by stanbon(75887)

Let amount of 11% solution be "x"
Then amount of 19% solution is "25-x" liters
Acid in 11% solution is 0.11x liters
Acid in 19% solution is 0.19(25-x)= 4.75-0.1925x liters
EQUATION:
acid + acid = 0.17(25) liters
0.11x+4.75-0.19x=4.25 liters
Multiply through by 100 to get:
11x+475-19x=425
-8x=-50
x=6.25 liters (amount of 11% solution)
25-x=18.75 liters (amount of 19% solution)
Cheers,
Stan H.